# Measuring Agents in Production

### v1

@@ Acknowledgements
We thank our many anonymous participants without whom beginning this study would not be possible. This research was supported by gifts from Accenture, Amazon, AMD, Anyscale, Broadcom Inc., Google, IBM, Intel, Intesa Sanpaolo, Lambda, Mibura Inc, Samsung SDS, and SAP. We also thank Alex Dimakis, Drew Breunig, Tian Xia, and Shishir Patil for their valuable feedback and insightful discussions.
Organization of Appendix
The appendix is organized as follows: in Appendix A we release the results on full survey dataset and report all statistics on the complete set of agent systems (not only the deployed subset used in the main paper), including comparisons between results on all agents and on deployed agents only; in Appendix B we provide additional analyses, such as our topic-normalization procedure for agent domains, the outlier domains where agents operate, and more detailed breakdowns of the challenges practitioners face when building agents. In Appendix C we unify our terminology for agents by defining what we mean by tasks, subtasks, and steps, and we provide an overview of the agent workflow schema used throughout the paper. In Appendix D we describe our interview protocol and present extended details of our case studies, including interview outlines and anonymized interviewee demographics. Finally, in Appendix E we provide the full survey questionnaire, including question wording and the branching logic that determines the survey flow.
Appendix A Results on All Survey Data
In the main body of the paper, we focused on results filtered exclusively to deployed agents in production or pilot phases, to highlight successful real-world practices under realistic operational constraints. In this appendix, we present the corresponding results computed over [All Data]: all 306 valid survey responses, regardless of deployment stage. This expanded view includes prototype, research, and legacy systems (sunset and retired), providing a broader perspective across the full agent development lifecycle.
For ease of comparison, each [All Data] figure mirrors a figure in the main text and uses the same layout and question wording. In the discussion below, we briefly summarize the key patterns in the full dataset and highlight how they compare to the deployed-only subset.
A.1 Applications, Users, and Requirements
Motivations and application domains.
Figure 14(a) reproduces our analysis of motivations for using agents over non-agentic alternatives using all survey responses. We observe that the overall ranking of benefits is highly stable compared to deployed only agents as shown in Figure 1: increasing productivity and efficiency remains the most frequently selected reason for adopting agents, followed by reducing human task-hours and automating routine labor .
For application domains of agents, (Figure 14(b)) become even more diverse in the full dataset compared to deployed only agents (Figure 2). The same high-level industries i.e., finance and banking, technology, and corporate services remain prominent. However, the long tail of “other” domains grows, reflecting additional experimental and research systems in areas such as education, creative tools, and scientific workflows that have not yet reached deployment.
End users and latency requirements.
Latency requirements also similarly remain relaxed in the full dataset (Figure 15(b) compared to Figure 5(b)). Most teams still report tolerating response times on the order of minutes, with a non-trivial fraction indicating that no explicit latency limit has been set. Compared to deployed agents only, the fraction of agents with undefined latency budgets is slightly higher in [All Data], which is consistent with prototypes and research artifacts that have not yet been hardened with production SLOs. Overall, these figures confirm that the preference for latency-relaxed, quality-focused applications is not an artifact of our deployment filter.
A.2 Models, Architectures, and Prompting
Core components and autonomy.
Figure 16 summarizes core component configurations and agent architectures over the full dataset and corresponds to Figure 8 in the main text. Compared to deployed agents (Figure 8(a)), distribution of number of distinct models used exhibits a heavier tail: non-deployed and research systems are more likely to combine many distinct models, leading to a higher incidence of configurations with four or more models. This pattern aligns with our qualitative observation that teams explore richer multi-model setups during early experimentation, then consolidate to a smaller, more manageable set of models as they move toward deployment.
Prompt lengths remain broadly similar between the deployed-only (Figure 8(b)) and full datasets ((Figure 16(b)) .
Comparison between Figure 8(c) and Figure 16(c) reveals a clearer separation in autonomy. When we include all agents, systems allow a greater number of autonomous steps or cycles before human intervention compared to deployed agents only. Experimental and research systems are more likely to sit in the “tens of steps” and “no explicit limit” regimes, whereas production deployments concentrate in the low-step regime to control cost, latency, and failure amplification. Taken together, these results reinforce our interpretation that bounded autonomy is a deliberate design choice for production reliability, while higher autonomy is more common in exploratory settings.
Prompt construction strategies and frameworks.
Figure 17(a) reproduces our analysis of prompt construction and framework usage using all survey responses. The results show that human-crafted prompts remain central across the full dataset: fully manual and manual+LLM strategies are still the dominant modes. However, when we include non-deployed agents, we see a modest shift toward more automated prompting: fully manual prompting is slightly more common among deployed agents (Figure 7), while the [All Data] distribution shows somewhat higher use of “fully autonomous” prompting and prompt optimizers. This suggests that automated prompt construction is currently used more as an experimental technique and is less frequently adopted in production systems where controllability is critical.
Figure 17(b) compares framework usage across all agents and corresponds to Figure 17(b) in the main text. The overall split between using any framework vs. no framework remains nearly identical between the full dataset and deployed-only subset, indicating that teams decide early whether to invest in a framework-based stack or implement their own orchestration. Within the “framework” group, the Other category expands slightly in [All Data], reflecting experimentation with a broader variety of less common or homegrown frameworks during research and prototyping stages, beyond the dominant framework families.
A.3 Evaluation Practices for All Agents
Figure 18 presents evaluation practices across all agents and mirrors Figure 10 in the main text. Figure 18(a) shows distribution of comparison against non-agentic baselines. When we include prototypes and research systems, the fraction of teams that explicitly compare their agents to alternative solutions is slightly lower than in the deployed-only subset (34% in all data vs. 38.7% in deployed agents in Figure 10(a)). This suggests that some experimental agents are perhaps still in early stages where rigorous baseline comparison has not yet been prioritized, or where teams are primarily exploring feasibility rather than relative gains.
Figure 18(b) reports the distribution of different evaluation methods. The ordering of methods remains unchanged: human-in-the-loop (manual) evaluation is still the most common strategy, followed by model-based evaluation (LLM-as-a-judge). However, manual evaluation is somewhat more prevalent among deployed agents (Figure 10(b)), whereas the [All Data] distribution shows a relatively higher share of automated methods. This is consistent with the idea that experimental and research systems may rely more on automated or lightweight checks, while production systems invest more heavily in human verification before and during deployment.
Figure 18(c) visualizes co-occurrence patterns between evaluation strategies. Human-in-the-loop evaluation remains the central hub in the evaluation graph, with high overlap with all other methods in the full dataset. At the same time, its co-occurrence with other strategies is slightly lower in [All Data] than in the deployed-only subset (Figure 10(c)), reflecting that some experimental systems use model-based or rule-based checks without consistently pairing them with human review. In contrast, deployed agents are more likely to combine automated evaluation with human verification perobably for higher assurance.
A.4 Data Handling, Latency Challenges, and Modalities Across All Agents
Figure 19 corresponds to Figure 12 and Figure 11 in the main text but reports statistics across all survey responses.
Figure 19 shows distribution of data sources for agents which is remarkably stable when moving from deployed agents only ( Figure 12) to [All Data] in Figure 19. This possibly indicates that the underlying data plumbing for agents is largely shared across lifecycle stages: teams tend to set up similar ingestion and handling pipelines during prototyping that then carry through to production with incremental hardening.
In addition, Figure 19(b) reports how often latency is described as a problem across all systems. The distribution changes only modestly when compared to the deployed-only subset (Figure 11), and we again see that latency is not the dominant blocker for most agent deployments. This supports our broader conclusion that agents are currently concentrated in latency-relaxed settings where quality and correctness dominate over strict real-time responsiveness.
Appendix B Analysis Details
B.1 Topic Normalization for RQ1
We then normalized these responses using an LLM-based semantic aggregation to merge semantically similar responses and categorize the answers for conceptual consistency and to avoid repeated or redundant categories. The resulting categories were manually reviewed to confirm semantic alignment. We then perform an LLM-based semantic mapping over each survey response to assign one or more domain labels. For example, the Healthcare Services category includes answers such as health care, medical, medicine, biomedicine, patient monitoring, virtual nursing, and care navigation. We use LOTUS [lotus] to perform the semantic aggregation and semantic mapping and provide the relevant program snippets below. We use gpt-4o-2024-08-06 as the LLM for both.
B.2 Outlier Domains
Table 2 presents the domains that were mentioned only once across all reported Agentic AI use cases. These outlier domains highlight the long-tail diversity of applications where Agentic AI is being explored beyond the dominant sectors such as finance, technology, and enterprise.
@@ B.2 Outlier Domains
B.3 Challenge Details
@@
We asked participants to rank the major categories of challenges they encounter during the development or operation of Agentic AI systems. Table 3 provides detailed descriptions of each challenge category identified in the survey, outlining the main technical and organizational issues practitioners reported when building Agentic AI systems. The five categories and their detailed definitions are provided in Table 3. Figure 21(b) illustrates how frequently each challenge category was assigned a given rank. For example, ‘Core Technical Performance’ was ranked as the most significant challenge (#1) by 16 respondents, by far more than any other category, indicating it remains the dominant source of difficulty in current Agentic AI system development. ‘Core Technical Performance’ encompasses a wide range of issues, including robustness, reliability, scalability, latency, and resource constraints. Its prevalence suggests that much of the community’s current effort is devoted to ensuring that systems perform consistently and dependably under real-world conditions. Following closely were ‘Data and Model Integrity’ and ‘System Integration and Validation’, both of which were reported as persistent sources of friction when transitioning systems from research prototypes to production environments. In contrast, ‘Transparency and Governance’ and Compliance and User Trust were ranked as lower-priority concerns. As shown in Figure 21(b), ‘Transparency and Governance’ was most frequently placed in the fifth position (14 occurrences), indicating that while practitioners recognize its long-term importance, it is not yet perceived as a primary bottleneck in current development cycles.
@@
Appendix C Terminology
Appendix D In-Depth Case Study Data
For building in-depth case studies, we arranged interviews with technical experts willing and able to supplement the published case study materials. Interviewers were selected to maintain organizational neutrality, assigned roles (lead, assistants, observers), and completed a series of pre-, post-, and in-interview procedures.
D.1 Interview Outline
The structure of interviews was determined by a preset list of 11 topic groups (below), and the availability of respective answers from open sources that need only be verified via interview. In general, interviewers were advised to prioritize topics 1–5 first, then 6–8, and finally 9–11 as time allowed.

### v2

@@ Acknowledgements
We thank our many anonymous participants without whom beginning this study would not be possible. This research was supported by gifts from Accenture, Amazon, AMD, Anyscale, Broadcom Inc., Google, IBM, Intel, Intesa Sanpaolo, Lambda, Mibura Inc, Samsung SDS, and SAP. We also thank Alex Dimakis, Aditya Parameswaran, Drew Breunig, Tian Xia, Bhavya Chopra, and Shishir Patil for their valuable feedback and insightful discussions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. All participant data was collected under confidentiality agreements and reported in aggregated, de-identified form. There may be other potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Organization of Appendix
The appendix is organized as follows. Appendix B provides supplementary results and deeper analyses that expand on the main findings, including extended breakdowns of agent requirements in RQ1 (e.g., users, domains, and application constraints) in Appendix B.1. Appendix B.2 expands on RQ2 results with the common architectural and engineering practices in production agents, including prompt construction strategies and recurring agentic workflow patterns. Appendices B.3 provides more detail on on evaluation practices that was discussed in RQ3. In addition, in Appendix B.4 we dig deeper into deployment challenges, with emphasis on security, privacy, and latency considerations.
Appendix C documents our study methodology, including the interview protocol and survey design, sampling, and data processing procedures. Appendix D reports results over the full survey dataset (beyond deployed agents only) and contrasts trends between deployed and non-deployed agents. Appendix E standardizes the terminology used throughout the paper (e.g., tasks, subtasks, and steps) and describes the workflow schema. Appendix F expands the related work discussion, situating MAP relative to prior surveys and empirical studies and clarifying how our study differs in scope and evidence. Finally, Appendix G contains the full survey instrument, including complete question text and branching logic.
Appendix B Supplementary Results and Analysis
In this appendix, we provide additional analysis and details related to the research questions discussed throughout the paper. We note that throughout the paper and appendix, for survey questions involving categorical comparisons, we report 95% confidence intervals computed using 1,000 bootstrap samples with replacement, where applicable.
B.1 Agents application Requirements
B.1.1 Agents Users
According to our study, the vast majority of agentic systems are designed to serve human users rather than other agents or software systems. Figure 8 shows the types and scale of users served by deployed agentic systems. As shown in this Figure, 92.5% of deployed agents report humans as their primary users. Among these, internal employees constitute the largest user group (52.2%), followed by external customers (40.3%). Only a small fraction of deployed systems (7.5%) primarily serve non-human consumers, such as downstream software or automated services.
Case study evidence suggests that the emphasis on internal users is often a deliberate deployment choice. Organizations frequently restrict early or initial deployments to internal settings to manage unresolved reliability, safety, and security risks. Internal deployments allow agent outputs to remain within organizational boundaries, where human oversight is readily available and errors typically carry lower external consequences. For example, several interviewed teams described internal operational agents that respond to employee requests, with human engineers able to intervene or override decisions when necessary. Across both internal- and external-facing systems, agents commonly support domain experts rather than operate autonomously. Many deployments require specialized domain knowledge to interpret agent outputs correctly—for instance, insurance authorization agents assisting nurses or incident response agents supporting site reliability engineers. This pattern reflects a broader role for agents as productivity-enhancing tools that augment expert workflows, with humans acting as final decision-makers or verifiers.
B.1.2 Application Domains
Figure 2 summarizes the application domains of deployed Agentic AI systems in our survey data, spanning 26 distinct domains. These domains reflect a broad range of real-world use cases, from enterprise and software operations to healthcare, scientific discovery, and communication services. In this section, we describe how domain labels were derived from free-text survey responses and how infrequently occurring domains (outliers) were handled.
Domain Normalization
The survey response for the question related to agent application domains (QN4 in Appendix Core Questions) was collected in free-text form. Therefore, we normalized the domain responses to enable systematic analysis. We first applied the state-of-the-art semantic parser LOTUS (Patel et al., 2025) to perform semantic aggregation, deriving a set of candidate domain categories by grouping semantically similar phrases (e.g., healthcare, medical, patient monitoring) into unified labels (e.g., Healthcare Services). Based on the LOTUS output and as shown in Figure 2, this resulted in nine categories in total: eight high-level application domains and one residual “Other” category capturing long-tail responses.
Each survey response was then independently annotated by three annotators, who assigned one or more relevant domain labels from this LOTUS-derived category set. We measured inter-annotator agreement across all pairwise combinations of the three annotators and across all labels, yielding a mean Cohen’s of 0.636. This level of agreement is commonly interpreted as substantial for multi-class semantic labeling tasks, particularly given the multi-select nature of the question and the fact that many deployed agent systems naturally span adjacent or overlapping domains.
Disagreements primarily arose in borderline cases where systems plausibly fit multiple related categories (e.g., Corporate Services vs. Finance & Banking, or Technology vs. Research & Development), rather than from ambiguity in category definitions. In cases where no consensus was reached among the three annotators, a fourth expert annotator adjudicated the final label. Figure 2 reports results based on these final consensus annotations. Additional implementation and normalization details, including representative LOTUS programs, are provided below in this appendix section.
@@ B.1.2 Application Domains
Outlier Domains
Figure 2 includes an “Other” category, which captures the long-tailed distribution of categories not shown explicitly in the figure. Table 1 presents domains that were mentioned only once across all reported deployed agent use cases. These outlier domains highlight the long-tail diversity of agent applications beyond the dominant sectors observed in the main analysis, illustrating how agentic systems are being explored across a diverse set of real-world problems.
B.1.3 Latency Requirements
Relaxed latency requirements are commonly observed among deployed agents. Figure 3 shows the distribution of maximum allowable end-to-end latency. Minutes is the most common target, followed by seconds. Notably, 17.0% report no defined limit yet. The latency tolerance reflects the productivity focus from Section 4.1. Agents are often used to automate tasks that can take humans hours to complete. Consequently, an agent taking multiple minutes to respond remains orders of magnitude faster than non-agentic baselines. Interview participants emphasized this advantage: even if an agent takes five minutes, that remains more than faster than assigning the task to a person on the team, especially when staffing shortages exist and the task is secondary to the user’s core responsibilities. Examples include nurses examining insurance details and software engineers responding to internal pager duty. Some deployed agents from case studies even batch requests hourly or overnight, further indicating latency is not a primary constraint.
However, this pattern breaks for real-time interactive applications. For example, practitioners building voice-driven systems report latency as their top challenge (Section B.4.2) during detailed case study. These systems compete against human conversation speeds rather than task completion baselines. Among our 20 detailed case studies, only 5 require real-time responsiveness. The remaining 15 cases tolerate extended processing times: 7 involve human review with relaxed timing, 5 operate as asynchronous background processes, and 3 have hybrid operation patterns. For these systems, processing times of minutes remain acceptable because the alternative can be days of human effort.
B.2 Agents Architecture
B.2.1 Agentic Frameworks
We find a divergence in framework adoption between survey respondents and interview case studies. As shown in Figure 9(a), among deployed agents from the survey, two-thirds (60.7%) use third-party agentic frameworks. Reliance concentrates around three primary frameworks: LangChain/LangGraph (30; LangChain Inc. (2025)) leads with 25.0%, followed by CrewAI (crewAI, 2025) at 10.7%, with LlamaIndex (LlamaIndex, 2025) and OpenAI Swarm (OpenAI, 2025b) both at 3.6%.
In sharp contrast, our detailed case studies reveal a strong preference for custom in-house agent implementations. Only 3 of 20 (15%) detailed case studies rely on external agent frameworks (2 use LangChain/LangGraph (30; LangChain Inc. (2025)), 1 uses DSPy (Khattab et al., 2024)). The remaining 17 teams (85%) build their agent application entirely in-house with direct model API calls. For example, one interview case explicitly shared that their agents are their own implementation of ReAct loops. Notably, two additional teams report starting with frameworks like CrewAI during the experimental prototyping phase but migrating to custom in-house solutions for production deployment to reduce dependency overhead.
B.2.2 Prompting
As shown in Figure 5 Across deployed systems, prompt construction remains predominantly human-driven, with limited adoption of fully automated methods. We further examine system prompt length distributions among our survey deployed agents in Figure 9(b). While a majority of agents use relatively concise prompts (51.5% under 500 tokens), prompt length exhibits a pronounced long tail. Among deployed agents, 24.2% use prompts between 500 and 2,500 tokens, 12.1% between 2,500 and 10,000 tokens, and an additional 12.1% exceed 10,000 tokens (Figure 9(b)). These longer prompts are rarely the result of a single design decision; instead, interview data suggests they often accumulate over time as systems incorporate guardrails, exception handling, policy constraints, and domain-specific instructions. As systems mature, prompts increasingly serve as centralized coordination artifacts rather than minimal task descriptions.
B.3 Evaluation
This section provides additional detail on how practitioners evaluate deployed agentic AI systems. We complement the results for RQ3 by exploring further on (i) whether deployed agents are explicitly compared against non-agentic alternatives, and (ii) the distribution of overlapping evaluation and verification strategies used in practice among deployed agents.
B.3.1 Comparison to baselines.
Figure 10(a) reports whether teams explicitly compare their deployed agentic systems against non-agentic baselines, such as existing software systems, traditional automated workflows, or human execution. 38.7% of respondents report conducting such comparisons, while the majority do not. In-depth interviews suggest several contributing factors: in some cases, agents are designed for tasks with no clear pre-existing alternative; in others, the baseline consists of a heterogeneous process involving multiple tools and human steps, making systematic technical comparison difficult. As a result, teams often evaluate agents based on task success, user outcomes, or qualitative improvements rather than direct performance deltas against a single baseline system.
B.3.2 Evaluation strategies overlap.
As discussed in § 6.2 and Figure 7, human-in-the-loop evaluation (74.2%) and LLM-as-a-judge approaches (51.6%) are the most commonly adopted evaluation strategies among deployed agentic systems. We further dive deeper into analyzing different evaluation strategies in Figure 10(b), where we explore which evaluation methods are commonly used together.
Figure 10(b) illustrates the co-occurrence of evaluation strategies across deployed agents, showing that human-in-the-loop evaluation is most frequently used in combination with other methods. Rather than relying on automated techniques in isolation, practitioners typically anchor rule-based verification, cross-referencing, and LLM-based judgments around human annotation. This pattern suggests that human judgment plays a central coordinating role in evaluation pipelines for production agentic systems.
This distribution indicates that many production agents operate in settings where correctness cannot be fully determined through deterministic rules or simple pattern matching. Instead, agents are deployed in domains requiring contextual understanding and nuanced judgment, such as customer support voice assistance or human resource operations. In these settings, practitioners rely on human and LLM-based evaluation to assess output quality, appropriateness, and task success, with automated checks serving a complementary role.
B.4 Challenges
@@
Like any emerging system, building agentic systems is not immune to challenges. We asked survey participants to rank the major categories of challenges they encounter during the development or operation of Agentic AI systems. Table 2 provides detailed descriptions of each challenge category identified in the survey, outlining the main technical and organizational issues practitioners reported when building Agentic AI systems. The five categories and their detailed definitions are provided in Table 2. Figure 11(b) illustrates how frequently each challenge category was assigned a given rank. For example, ‘Core Technical Performance’ was ranked as the most significant challenge (#1) by 37.9% of respondents, by far more than any other category, indicating it remains the dominant source of difficulty in current Agentic AI system development. ‘Core Technical Performance’ encompasses a wide range of issues, including robustness, reliability, scalability, latency, and resource constraints. Its prevalence suggests that much of the community’s current effort is devoted to ensuring that systems perform consistently and dependably under real-world conditions. Following closely as shown in Figure 11(a) ‘Data and Model Integrity’ and ‘System Integration and Validation’ were reported as second-ranked persistent sources of friction when transitioning systems from research prototypes to production environments. In contrast, ‘Transparency and Governance’ and Compliance and User Trust were ranked as lower-priority concerns, indicating that while practitioners recognize its long-term importance, it is not yet perceived as a primary bottleneck in current development cycles.
@@
B.4.1 Security and Privacy Challenges
Security and privacy consistently rank as secondary concerns in both of our studies, with practitioners prioritizing output quality and correctness. Figure 11(a) shows that Compliance and User Trust ranks fourth among challenge categories. Given that §4.3 shows 52.2% of systems serve internal employees and many systems with human supervision, this prioritization reflects current deployment environments and requirements rather than dismissing security’s importance.
Data ingestion and handling. Survey results in Figure 12(a) show that 89.7% of systems ingest information from databases, 65.5% ingest real-time user input, and 51.7% ingest other real-time signals. Notably, 69.0% of systems retrieve confidential or sensitive data, while only 34.5% retrieve persistent public data. Given the high prevalence of sensitive data usage and user inputs, preserving privacy is critical. Our interview case studies reveal that teams address this through legal methods. For example, a team building healthcare agents report relying on standard data-handling practices and strict contractual agreements with model providers to prevent training on their user data.
Security practices. In-depth interview participants describe four approaches to managing security risks through constrained agent design. First, six teams restrict agents to “read-only” operations to prevent state modification. For example, one SRE agent case study generate bug reports and proposes action plans, but leaves the final execution to human engineers. Second, three teams deploy agents in sandboxed or simulated environments to isolate live systems. In one instance, a code migration agent generates and tests changes in a mirrored sandbox, merging code only after software verification. Third, one team builds an abstraction layer between agents and production environments. This team constructs wrapper APIs around production tools, restricting the agent to this intermediate layer and hiding internal function details. Finally, one team enforces role-based access controls that mirror agent user’s permissions. However, the agent team reports this remains challenging, as agents can bypass these configurations when accessing tools or documents with conflicting permissions.
B.4.2 Latency Challenges
We examine the degree to which agent execution latency hinders deployment. Survey results indicate that latency represents a manageable friction rather than a hard stop for most teams. Figure 12(b) shows that only 14.8% of deployed survey agents identify latency as a critical deployment blocker requiring immediate resolution, while the majority (59.3%) report it as a marginal issue, where current latency is suboptimal but sufficient for deployment. We suspect that this tolerance correlates with the prevalence (15/20 in detailed interview case studies) of asynchronous agent execution paradigm (Section B.1.3) and (52.2% from survey) internal user bases (Section 4.3). Notably, we observe a consistent latency distribution across the full survey dataset, including experimental systems (Figure 24(b)). We believe this consistency signals a broader preference for building offline agents, as discussed in Section B.1.3.
Interactive agent latency requirements. While latency is not a critical challenge for most agent applications, it remains a critical bottleneck for real-time interactive agents. Two interviewed teams, building voice agents, report continuous engineering efforts to match human conversational speeds. Unlike asynchronous workflows, these systems require seamless turn-taking where delays disrupt the user experience. Achieving fluid real-time responsiveness beyond rigid turn-based exchanges remains an open research question and development challenge.
B.4.3 Modalities
Supporting multiple data modalities emerges as an additional deployment challenge for production agent systems. Beyond text-based interactions, practitioners increasingly aim to extend agents to handle richer inputs and outputs, including speech, images, video, spatiotemporal data, and domain-specific scientific formats. While these capabilities unlock new application opportunities, they also introduce substantial engineering, evaluation, and reliability challenges.
To understand current practice and future direction, we asked survey participants to report (1) which data modalities their deployed agents currently support and (2) which additional modalities they plan to add. Figure 13 summarizes these responses for deployed agent systems, contrasting modalities already supported (red) with those planned for future support (blue). As shown in this Figure, text dominate current deployments: 93% of surveyed production agents accept natural-language text input. In contrast, support for non-textual modalities remains relatively limited today. However, the strongest expected growth is concentrated precisely in these less mature modalities. Bars to the left of the dashed line indicate modalities with anticipated increases in future support, while modalities to the right are already relatively more adopted with comparatively limited planned expansion. Notably, the modalities with the largest projected growth—such as images, video, spatiotemporal data, and scientific data—are all non-textual, pointing toward increasingly multimodal agent systems.
Interview data helps contextualize this pattern. Participants consistently report that early deployments prioritize modalities that are easy to measure, validate, and debug. Text-based agents benefit from established evaluation workflows, human verification, and relatively fast iteration cycles. In contrast, multimodal agents face weaker correctness signals, higher infrastructure complexity, and more expensive data pipelines. As a result, teams often defer multimodal expansion until text-centric deployments achieve sufficient reliability.
This trend presents both challenges and opportunities. On the challenge side, multimodal agents complicate evaluation, observability, and failure detection, exacerbating issues already present in text-only systems. On the opportunity side, growing practitioner demand suggests a clear need for research on multimodal agent architectures, modality-aware evaluation methods, and system designs that support reliable multimodal execution at runtime. As production agents mature, we expect modality expansion to follow successful stabilization of text-based deployments, extending agent capabilities toward richer, more complex real-world inputs.
Appendix C Collected Data Details
In this section, we provide additional details about our collected data sources, including (i) the survey respondent population and the deployment characteristics of the reported agentic systems (Appendix C.1), and (ii) supplementary information on the in-depth interview case studies such as demographics and organizational context (Appendix C.2). In total, our survey collected 306 valid responses and, together with 20 in-depth interviews, this data forms the empirical basis of this study on AI agents in production.
| Business Operations |
| C01: Insurance claims workflow automation |
| C02: Customer care internal operations assistance |
| C03: Human resources workflow automation and assistance |
| Communication Tech, Multi-lingual Multi-dialect |
| C04: Communication automation services |
| C05: Automotive communication services |
| Scientific Discovery |
| C06: Biomedical sciences workflow automation |
| C07: Materials safety and regulatory analysis automation |
| C08: Chemical data interactive exploration |
| Software DevOps |
| C09: Spark version code and runtime migration |
| C10: Software development life cycle assistance end-to-end |
| C11: Software engineer/developer slack support |
| C12: SQL optimization |
| C13: Code auto-completion and syntax error correction |
| Software & Business Operations |
| C14: Data analysis and visualization |
| C15: Enterprise cloud engineer and business assistance |
| C16: Site reliability incident diagnoses and resolution |
| C17: Software products technical question answering |
C.1 Survey Data Details
Survey respondents self-identified as practitioners actively building AI agents across 26 application domains (Figure 2), spanning areas such as finance, healthcare, and legal services. Of the 306 respondents, 294 reported that they have directly contributed to building and designing at least one agent system. As reported in Figure 14(a) (), among those who disclosed their role, respondents are predominantly technical professionals, with a large share identifying as software and machine learning engineers. Figure 14(b) reports the deployment stages of the agentic systems respondents contributed to (); among those who reported deployment stage, 82% indicated that their systems are in production or pilot phases, reflecting rapid transition from experimental prototypes to real-world deployments.
Beyond deployment stage, we examine the scale of the user base for deployed systems. Figure 14(c) summarizes reported end-user counts (), showing substantial variation in deployment scale. In particular, 42.9% of reported deployments serve user bases in the hundreds, while 25.7% serve tens of thousands to over one million daily users. Together, these distributions indicate that our survey captures both smaller deployments and high-impact systems operating at significant scale, motivating our focus on deployed agents in the main analysis.
C.2 In-Depth Case Study Details
This appendix provides details on the in-depth case studies to support the qualitative findings. In total, we curated 20 interview-based case studies, selected to reflect diversity in application settings, organizational maturity, and geographic reach. The anonymized case studies and their representative use-case descriptions are summarized in Table 3. These cases span multiple application categories, including business operations (C01-C03), communication technologies and multilingual or multidialect systems (C04-C05), scientific discovery (C06-C08), software DevOps and infrastructure (C09-C13), and software and business operations (C14-C16). Each case is referenced throughout the paper using anonymized identifiers (C01, C02, …). When organizations operated multiple agent deployments, we prioritized selecting distinct use cases to avoid over-representing any single institution and to capture a broader range of deployment patterns.
C.2.1 Deployment Stage and Organizational Maturity
All interviewed systems serve real-world users: 14 cases are in full production and 6 are in final pilot phases. The studied systems support both internal users (5 cases) and external enterprise users (15 cases), and originate from organizations spanning a wide range of maturity levels—from seed-stage startups to large enterprises with global footprints (Figure 15). To respect confidentiality agreements with case study sources, we report only aggregate statistics about organizational characteristics and geographic footprint.
C.2.2 Geographic distribution.
Figures 17 and 17 summarize the geographic footprint of organizations in our case studies. As shown in Figure 17, case study sources operate across a wide range of country-level presence, from organizations active in a single country to globally distributed deployments spanning hundreds of countries. Figure 17 shows that case study sources operate across one to six continents, indicating that the interviewed systems include both regionally focused deployments and globally distributed services. These distributions suggest that the qualitative findings reflect agent deployments operating under heterogeneous regulatory, linguistic, and operational environments, rather than being confined to a single geographic context.
C.2.3 Participant Background, Recruitment, and Data Collection Timeline
We conducted interviews with technical practitioners directly responsible for the design, implementation, or operation of the studied agentic systems, including ML engineers, software engineers, and senior technical leads. To protect confidentiality, we do not report individual-level demographic attributes; however, participants represented a range of engineering roles and experience levels spanning system architecture, deployment, evaluation, and ongoing operations. Although studied systems may have been publicly announced, public-facing materials (e.g., product documentation, marketing releases, or high-level technical overviews) do not capture the implementation- and operations-level detail central to our analysis. As such, access to practitioners with the required depth of expertise typically occurs through professional networks and practitioner-oriented events. Participants were therefore recruited via the authors’ professional networks as well as outreach through presentations at agent-focused technical venues (anonymized), and were screened to ensure active, hands-on involvement in the development or maintenance of the systems under study.
Data collection occurred across multiple phases: initial survey and interview design began in March, followed by interviews with upstream stakeholders from mid‑April through mid‑May. Three subsequent recruitment rounds were conducted through agent-focused technical events spaced roughly two months apart—late May, early August, and mid‑October—with the final interview completed in November.
Considerations for fast‑evolving systems. Because agentic systems and their supporting infrastructure evolve rapidly, collecting data across multiple phases allowed us to capture changes in engineering practices as they emerged over the study period. This staggered timeline also broadened the range of contexts represented, as participants engaged with different system states, model versions, or operational conditions. However, we recognize that later interviews may reflect different technical environments than earlier ones, which can complicate synthesis and introduce recency‑related bias. To mitigate these issues, we analyzed interviews with close attention to their temporal context, distinguishing themes that appeared consistently across phases from those tied to time‑specific system developments.
C.3 Interviews Details
C.3.1 Standardized Interview Procedures.
We followed a consistent sequence of pre-, in-, and post-interview procedures across all case studies. Interviews adhered to a shared semi-structured protocol spanning 11 topic areas (Appendix C.3.2), covering system architecture, evaluation practices, deployment challenges, operational constraints, and measures of agent value. Each interview lasted 30–90 minutes and involved 2–5 participants with fixed roles to improve consistency and reduce interviewer effects. Interviewers were selected to maintain organizational neutrality with respect to participating teams.
Pre-interview context gathering, when applicable, was restricted to publicly available sources (e.g., product documentation or engineering blogs) to avoid introducing private or leading assumptions. Depending on participant preference, interviews were either recorded or documented through detailed human notes. Post-interview summaries were cross-validated among interviewers to ensure accuracy and internal consistency. In accordance with confidentiality agreements, all data are anonymized and reported only in aggregate. Interview discussions were guided by the predefined topic groups, with interviewers instructed to prioritize topics 1–5, followed by 6–8, and then 9–11 as time permitted. Topics with available public information were used primarily for verification rather than open-ended elicitation.
C.3.2 Interview Outline